OpenStudy (anonymous):

Need help! It seems like an easy question, but I can't figure it out. Marvel Man movie is being released to a limited number of movie theaters. The number of showtimes for the movie is twice the number of movie theaters. The number of seats available for each showtime is three times the number of showtimes. Let x represent the number of theaters showing the movie. a.Determine the three factors that are to be multiplied to determine the total number of seats available for the movie. b.When simplified what is the degree and classification of this expression?

OpenStudy (anonymous):

The factors would be x 2x and 6x?

jimthompson5910 (jim_thompson5910):

Let's declare the following x = # of theaters y = # of showtimes z = # of seats for each showtime (assume each theater has the same number of seats in the same sized room)

jimthompson5910 (jim_thompson5910):

" The number of seats available for each showtime is three times the number of showtimes." we are told that "The number of showtimes for the movie is twice the number of movie theaters." so we know (# of showtimes) = 2*(# of theaters) y = 2x

jimthompson5910 (jim_thompson5910):

also we know " The number of seats available for each showtime is three times the number of showtimes." so (# of seats per showtime) = 3*(# of showtimes) z = 3y z = 3(2x) z = 6x

jimthompson5910 (jim_thompson5910):

So I agree with your answers of x, 2x and 6x You multiply x,y,z (aka x,2x,6x) out to get the total number of seats overall

jimthompson5910 (jim_thompson5910):

what did you get for part b?

OpenStudy (anonymous):

I got 12x^3??? So it would be a third degree trinomial?

jimthompson5910 (jim_thompson5910):

it's 3rd degree, but it's not a trinomial

jimthompson5910 (jim_thompson5910):

a trinomial has 3 terms ex: 3x^2 + 5x + 6

OpenStudy (anonymous):

sorry didnt mean trinomial. I meant monomial

jimthompson5910 (jim_thompson5910):

yep, 3rd degree monomial or cubic monomial

OpenStudy (anonymous):

Thank you. There's one more part. c. Interpret each pairing as a single unit in the context of the problem? I dont understand what its wanting

jimthompson5910 (jim_thompson5910):

each pairing of factors?

OpenStudy (anonymous):

yeah sorry.

jimthompson5910 (jim_thompson5910):

ok it might help to make a 3x3 table like this 

jimthompson5910 (jim_thompson5910):

then along the top and left side, we'll have x, 2x, 6x